OpenStudy (anonymous):

The company discovered that it costs $30 to produce 2 widgets, $118 to produce 4 widgets, and $766 to produce 10. Find the cost of producing 6 widgets. Use the quadratic function to solve

OpenStudy (anonymous):

You're told the function you should be looking for is quadratic, and you have three (x,y) points to use. The general form of the quadratic is\[y=ax^2+bx+c\]If you substitute each pair of points, you'll end up with three linear equations in the constants, a, b and c. You can then solve these using linear algebra. So, for the first point, (2,30),\[30 = a(2^2)+b(2)+c \rightarrow 30=4a+2b+c\]That's equation 1.

OpenStudy (anonymous):

For (4,118),\[118=16a+4b+c\]and for (10,766),\[766=100a^2+10b+c\]

OpenStudy (anonymous):

sorry, typo on last one. Should be\[766=100a+10b+c\]

OpenStudy (anonymous):

Can you take it from here? You use these three equations to solve for a, b and c. Then you can set up your quadratic, and plug in x=6 to get the cost.

OpenStudy (anonymous):

probably not. seeing i'm completely confused and my online courses don't really do enough explaining for math.